Simplify:
-3 sqrt50
sqrt(18ab) * sqrt(6ab^3)
sqrt(54)/sqrt(24)
sqrt(1/7)
sqrt(12) + 2(sqrt27) - sqrt 48

Answer by checkley77(12569)

-3 sqrt50
-3sqrt(25*2)=-3*5sqrt2=-15sqrt2
------------------------------------
sqrt(18ab) * sqrt(6ab^3)
sqrt(18*6a^2b^4)=sqrt(3*6*6a^2b^4)=6ab^2sqrt(3)
-------------------------------------------------
sqrt(54)/sqrt(24)=sqrt(9*6)/sqrt(4*6)=3sqrt(6)/2sqrt(6)=3/2=1.5
--------------------------------------------------------------------
sqrt(1/7)=sqrt.1428=.378
----------------------------
sqrt(12)+2(sqrt27)-sqrt(48)=sqrt(4*3)+2sqrt(9*3)-sqrt(16*3)=2sqt3+2*3sqrt3-4sqrt3=2sqrt3+6sqrt3-4sqrt3=4sqrt3=6.9282